OpenStudy (anonymous):

An artist wants to make alabaster pyramids using a block of alabaster with a volume of 576 cubic inches. She plans to make each pyramid with a square base area of 3 square inches and a height of 4 inches. At most, how many pyramids can the artist make from the block of alabaster? 136 144 154 166 172 HELP!!!!

OpenStudy (anonymous):

So what you would do is V = 1/3 * area of base * height V = 1/3 * 3 * 4 V = 4 in^3 576/4 = 144 => max number of pyramids made out of the block of alabaster.

OpenStudy (anonymous):

Does that kinda help?

OpenStudy (anonymous):

Yes thank you so much!!

OpenStudy (anonymous):

Yea No problem if you need anything don't hesitate to ask. XD